OpenStudy (anonymous):

Find the zeros of the function h(x) = x2 +4x by factoring. What is the smallest zero of the function? Enter a number only.

OpenStudy (anonymous):

\[\large h(x)=x^2+4x=x(x+4)=0\] remember the zero product property

OpenStudy (anonymous):

i don't get it would it be x = 0 and x = -4

OpenStudy (anonymous):

yes x=0 and x=-4

OpenStudy (anonymous):

ok thanks :p